OpenStudy (anonymous):

Verify the identity: [sin(3t)-sin(t)]/[cos(3t)+cos(t)]=tan(t)

OpenStudy (anonymous):

For sin: sin(3t)*cos(t)-cos(3t)*sin(t) For cos: cos(3t)*cos(t)-sin(3t)*sin(t) Considering that, I want to say that this is not an identity.

OpenStudy (anonymous):

sin3t=sin(2t+t)=sin2tcost+cos2tsint cos3t=cos(2t+t)=cos2tcost-sin2tsint

OpenStudy (anonymous):

plug them in the left side of the equation, you will get:

OpenStudy (anonymous):